OpenStudy (anonymous):

Medal Write an equivalent fraction with the given denominator. 2/ 9 = ?/36

OpenStudy (hockeychick23):

2/9=8/36

OpenStudy (anonymous):

how is the answer 8

OpenStudy (sleepyjess):

Lets set up equivalent fractions. \(\dfrac{2}{9} = \dfrac{x}{36}\) Do you know how to cross multiply?

OpenStudy (hockeychick23):

you use cross multiplication to get the answer. So you do 2*36 and 9*x and set them equal to each other

OpenStudy (anonymous):

yes

OpenStudy (sleepyjess):

Ok, so what do you get after cross multiplying?

OpenStudy (anonymous):

2 *36 = 72

OpenStudy (sleepyjess):

Yep, and 72 = 9x. Divide both sides by 9

OpenStudy (anonymous):

72 divide 8 =9

OpenStudy (sleepyjess):

72/9, not 8 :)

OpenStudy (sleepyjess):

but is 72/8 = 9, then 72/9 = 8 :)

OpenStudy (sleepyjess):

*but if

OpenStudy (anonymous):

so the answer is 8?

OpenStudy (anonymous):

yes :)

OpenStudy (sleepyjess):

Yep ^_^

OpenStudy (sleepyjess):

\(\dfrac{2}9 = \dfrac8{36}\)

OpenStudy (anonymous):

:) thank u everyone

OpenStudy (sleepyjess):

No problem! :) Thank you for not just taking the answer, but instead asking, how?

OpenStudy (sleepyjess):

I applaud you for that. There are not many that will do that.